Lily, Mex and Kim spent totally $34 on the ice cream shop. Lily spent $2 more than Kim spent, Kim spent two third the amount of money that Mex spent. How much money did each of them spend?
Please give me some idea to solve this problem. Thanks!

Answer by Cromlix(4381)

Hi there,
See what you think of this:-
Make Mex spend x
Kim spent 2/3x
Lily spent 2 + 2/3x
Mex + Kim + Lily = $34
x + 2/3x + 2 + 2/3x = $34
Collect like terms
x + 2/3x + 2/3x = 34 - 2
x + 4/3x = 32
7/3x = 32
x = (32 * 3)/7 (* times)
x = $13.71 (2 decimal places)
Mex spent $13.71
Kim spent $9.14
Lily spent $11.14
$13.71 + $9.14 + $11.14 = $33.99
Hope this helps :-)